How much work is done on a 5 kg block if it is to be moved on a ramp 30° above the horizontal and height of 1 meter? What are the values of work done by each force?
Answer #1

Force acting on the block = weight = mg

component of weight along the slope = mg sin a (downwards) = 5*9.8*sin 30 = 24.5 N

displacement = 1/ sin 30 = 2 m

so work done = F*S cos a

where F is the force , S = displacement and a = angle between force and dispalcement

W = 24.5*2* cos 180 = -49 J

there is only one force acting and that is weight

work done = FS cos a = mg S* cos a

= 5*9.8*2*cos (90 + 30)

= -49 J